Convert to a logarithmic equation.
Show work please
e^2=7.3891

Answer by Alan3354(69443)

Convert to a logarithmic equation.
Show work please
e^2=7.3891
==============
ln(e^2) = ln(7.3891)
2 = ln(7.3891)